# Quantum Gravitational Corrections in Cosmology

### v1

@@
Xavier Calmet111x.calmet@sussex.ac.uka, Roberto Casadio222casadio@bo.infn.itb,c and Marco Sebastianutti333m.sebastianutti@sussex.ac.uka
aDepartment of Physics and Astronomy, University of Sussex, Brighton, BN1 9QH, United Kingdom
bDipartimento di Fisica e Astronomia, Università di Bologna, via Irnerio 46, 40126 Bologna, Italy
cI.N.F.N., Sezione di Bologna, IS - FLAG, via B. Pichat 6/2, 40127 Bologna, Italy
dAlma Mater Research Center on Applied Mathematics - AM2, via Saragozza 8, 40123 Bologna, Italy
@@ 1 Introduction
In this paper we use the unique effective action of quantum gravity [1, 2, 3, 4, 5, 6] to calculate quantum gravitational corrections to several cosmology models. There have been previous studies of quantum gravitational effects in cosmology using the concept of effective action but these studies have limited themselves to either the local part of the effective action [7] or its non-local part [8, 9, 10, 7] and these results are thus inconsistent as they are not renormalization group invariant. In [11], the collapse of a dust ball has been considered using the Oppenheimer-Snyder model. The interior metric is given by the Friedmann-Lemaître-Robertson-Walker (FLRW) metric. This paper obtained for the first time quantum gravitational corrections to the FLRW metric which are renormalization group invariant. Note that in this model the classical singularity is in the future and these results are thus not directly applicable to standard cosmological models. The aim of our new paper is to extend these results to cosmology to demonstrate how to calculate reliably quantum gravitational corrections to cosmological models and to discuss their applicability domain.
@@ 3 Quantum gravitational corrections to FLRW cosmology
$\hat{c}_{1}(\mu)=\hat{c}_{1}(\mu_{\star})-2\hat{\alpha}\log{\left(\frac{\mu}{\mu_{\star}}\right)}.$ (30)
@@
$H(t)=\frac{2}{3t}+\frac{32\pi{t_{\rm P}}^{2}\hat{c}_{1}(\mu_{\star})}{t^{3}}+\frac{64\pi{t_{\rm P}}^{2}\hat{\alpha}}{t^{3}}\left[-\frac{2}{3}+\log{\left(-\mu_{{\rm E}\star}t\right)}\right]+{\cal O}(H_{q}^{2}).$ (32)
@@
$a(t)=a_{0}\left(\frac{t}{t_{0}}\right)^{\frac{2}{3}}\bigg{\{}1-\frac{16\pi{t_{\rm P}}^{2}\hat{c}_{1}(\mu_{\star})}{t^{2}}+\frac{16\pi{t_{\rm P}}^{2}\hat{\alpha}}{t^{2}}\left[\frac{1}{3}-2\log{\left(-\mu_{{\rm E}\star}t\right)}\right]\bigg{\}}+C_{1}t^{\frac{2}{3}}+{\cal O}(a_{q}^{2}),$ (33)
@@
We now plot the scale factors for an expanding, Fig. 1, and collapsing, Fig. 2, dust-filled flat FLRW universe. We vary $\hat{c}_{1}(\mu_{\star})$ and fix the parameters: $\mu_{\rm E}^{\star}=\mu_{\star}\cdot e^{\gamma_{\rm E}}\sim 1.78\mu_{\star}$, $C_{1}=\xi=0$, $N_{S}=4$, $N_{G}=1$, which are the Standard Model values, $(\mu_{\star})^{-1}=M_{P}\sqrt{160\pi/N}$, $a_{0}=1$ and ${t_{\rm P}}=\pm 1$, $t_{0}=\pm 100$, for the expanding and contracting case respectively. Note that times are in Planck time units.
@@ 4 Conclusions
In this paper we have shown how to reliably calculate quantum gravitational corrections to cosmological models. We have shown that it is important to consider the full effective action to obtain renormalization group invariant solutions. We have investigated the validity range of our techniques within simple cosmological models. Any large effect such as bounces cannot be described reliably within our approximations as by definition in such cases, the quantum corrections dominate over the classical solution. However, our results can be considered as tantalising hints that quantum corrections could indeed lead to bounces in the late-time cosmological models and accelerated expansion in the past of the universe.

### v2

@@
[…]
1Department of Physics and Astronomy, University of Sussex, Brighton, BN1 9QH, United Kingdom
2Dipartimento di Fisica e Astronomia, Università di Bologna, via Irnerio 46, 40126 Bologna, Italy
3I.N.F.N., Sezione di Bologna, IS - FLAG, via B. Pichat 6/2, 40127 Bologna, Italy
4Alma Mater Research Center on Applied Mathematics - AM2, via Saragozza 8, 40123 Bologna, Italy
@@ 1 Introduction
In this paper we use the unique effective action of quantum gravity [1, 2, 3, 4, 5, 6] to calculate quantum gravitational corrections to several cosmology models. [^1] The derivation of an effective action involves the splitting of the metric $g=\bar{g}+h$ in to a background $\bar{g}$ and a perturbation $h$. The perturbation $h$ is then integrated out. To perform this $h$-integration, one must fix a gauge. In general, this gauge choice will also fix the gauge for $\bar{g}$ and it thus results in an effective action that is gauge dependent. This problem can be resolved using DeWitt’s method of mean-field gauges. A further complication that arises is that the off-shell effective action in gauge theories depends on the choice of parameterization of the quantum fields. A solution to this issue was proposed in [7], leading to the unique effective action. For gravity, this effective action is constructed using covariant perturbation theory or the generalized Schwinger-DeWitt technique. There have been previous studies of quantum gravitational effects in cosmology using the concept of effective action but these studies have limited themselves to either the local part of the effective action [8] or its non-local part [9, 10, 11, 8] and these results are thus inconsistent as they are not renormalization group invariant. In [12], the collapse of a dust ball has been considered using the Oppenheimer-Snyder model. The interior metric is given by the Friedmann-Lemaître-Robertson-Walker (FLRW) metric. This paper obtained for the first time quantum gravitational corrections to the FLRW metric which are renormalization group invariant. Note that in this model the classical singularity is in the future and these results are thus not directly applicable to standard cosmological models. The aim of our new paper is to extend these results to cosmology to demonstrate how to calculate reliably quantum gravitational corrections to cosmological models and to discuss their applicability domain.
@@ 3 Quantum gravitational corrections to FLRW cosmology
$\hat{c}_{1}(\mu)=\hat{c}_{1}(\mu_{*})-2\hat{\alpha}\log{\left(\frac{\mu}{\mu_{*}}\right)}.$ (30)
@@
$H(t)=\frac{2}{3t}+\frac{32\pi{t_{\rm P}}^{2}\hat{c}_{1}(\mu_{*})}{t^{3}}+\frac{64\pi{t_{\rm P}}^{2}\hat{\alpha}}{t^{3}}\left[-\frac{2}{3}+\log{\left(-\mu_{\rm E}^{*}t\right)}\right]+{\cal O}(H_{q}^{2}).$ (32)
@@
$a(t)=a_{0}\left(\frac{t}{t_{0}}\right)^{\frac{2}{3}}\bigg{\{}1-\frac{16\pi{t_{\rm P}}^{2}\hat{c}_{1}(\mu_{*})}{t^{2}}+\frac{16\pi{t_{\rm P}}^{2}\hat{\alpha}}{t^{2}}\left[\frac{1}{3}-2\log{\left(-\mu_{\rm E}^{*}t\right)}\right]\bigg{\}}+C_{1}t^{\frac{2}{3}}+{\cal O}(a_{q}^{2}),$ (33)
@@
We now plot the scale factors for an expanding, Fig. 1, and collapsing, Fig. 2, dust-filled flat FLRW universe. We vary $\hat{c}_{1}(\mu_{*})$ and fix the parameters: $\mu_{\rm E}^{*}=\mu_{*}\cdot e^{\gamma_{\rm E}}\sim 1.78\mu_{*}$, $C_{1}=\xi=0$, $N_{S}=4$, $N_{G}=1$, which are the Standard Model values, $(\mu_{*})^{-1}=M_{P}\sqrt{160\pi/N}$, $a_{0}=1$ and ${t_{\rm P}}=\pm 1$, $t_{0}=\pm 100$, for the expanding and contracting case respectively. Note that times are in Planck time units.
@@ 4 Conclusions
In this paper we have shown how to reliably calculate quantum gravitational corrections to cosmological models. We have shown that it is important to consider the full effective action to obtain renormalization group invariant solutions. We have investigated the validity range of our techniques within simple cosmological models. Any large effect such as bounces cannot be described reliably within our approximations as by definition in such cases, the quantum corrections dominate over the classical solution. However, our results can be considered as tantalising hints that quantum corrections could indeed lead to bounces in the late-time cosmological models and accelerated expansion in the past of the universe. Note that our work could easily be extended to a multi-component cosmological model or even models with a general perfect fluid with a variable equation of state. However, given the smallness of the effect, there is little value in doing actual phenomenology. The main contribution of our work is to show that quantum gravitational corrections in cosmology are calculable from first principles.